Given events J and K: P(J) = .18; P(K) = .37; P(J OR K) = .45
Find P(J AND K).

Find the probability of the complement of event (J AND K).

Find the probability of the complement of event (J OR K).

Answer by ikleyn(53751)

.
Given events J and K: P(J) = .18; P(K) = .37; P(J OR K) = .45
Find P(J AND K).
Find the probability of the complement of event (J AND K).
Find the probability of the complement of event (J OR K).
~~~~~~~~~~~~~~~~~~~~~~

Start from this probability IDENTITY, which is valid for all/every/each events 

    P(J or K) = P(J) + P(K) - P(J and K).


Substitute here the given values.  You will get 

    0.45 = 0.18 + 0.37 - P(J and K).


From this equation, find  P(J and K)

    P(J and K) = 0.18 + 0.37 - 0.45 = 0.1.


ANSWER.  P(J and K) = 0.1.

Solved.


To get the complementary probability, subtract the base probability from 1.